Session: Feb 1999 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (D. Dallacasa at Tue Feb 23 08:06:00 MET 1999) Success
The observation went smoothly. Sky clear. Some interference at low elevation, but it should not be a problem

Torun (K.Borkowski at Tue Feb 23 09:33:00 MET 1999) Success
No known problems.

Torun (K.Borkowski at Tue Feb 23 09:38:04 MET 1999) Success
No known problems.

Noto (C.Trigilio at Tue Feb 23 15:50:43 MET 1999) Success
Strong wind, some interference.

Onsala (A. Polatidis at Tue Feb 23 18:53:50 MET 1999) Success
WE are observing using the backup maser. The formatter was re-syncrhonised before the start of the experiment.

Effelsberg (dgraham at Wed Feb 24 09:43:09 MET 1999) Failure
whole experiment zapped by snow. no tape sent.

Jodrell (Alastair Gunn at Wed Feb 24 12:50:16 MET 1999) Minor failures
FS restarted at 21:32 to set correct offsets. Cal temps used were wrong in procedure file, should be 15.2 and 15.6 K. No other known problems.

Westerbork (Tony Foley at Wed Feb 24 19:50:18 MET 1999) Success
This waz the first THIN tape observation after changing the write stack slope. Data looked OK, except for the large offset of tracks on tape.

Calibration (Jean\-Francois Desmurs at Wed Mar 3 11:20:40 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb99 directory under the name: el023.antab. Read the comments in the file to get further information. (note: Most Ef Tsys values wrong)